Vail Book Mart sells books and other supplies to students in a state where the sales tax rate is 7 percent. Vail Book Mart engaged in the following transactions for 2018. Sales tax of 7 percent is collected on all sales. 1. Book sales, not including sales tax, for 2018 amounted to $270,000 cash. 2. Cash sales of miscellaneous items in 2018 were $151,000, not including tax. 3. Cost of goods sold was $209,000 for the year. 4. Paid $131,000 in operating expenses for the year. 5. Paid the sales tax collected to the state agency.

a. What is the total amount of sales tax Vail Book Mart collected and paid for the year?

b. What is the Vail Book Mart's net income for the year?
Answer #1

ANSWER

a)    Amount of sales tax collected and paid = 270,000 *0.07+151,000*0.07 =18900 + 10570 $ 29,470
b) Net Income:
Total sales revenue 421,000
Cost of goods sold 209,000
Operating expenses 131,000
Net Income 81,000

_____________________________________________

If you have any query or any Explanation please ask me in the comment box, i am here to helps you.please give me positive rating.

*****************THANK YOU**************